Pepe, Incorporated acquired 60% of Devin Company on January 1, 2017. On that date Devin sold equipment to Pepe for $45,000. The equipment had a cost of $120,000 and accumulated depreciation of $66,000 with a remaining life of 9 years. Devin reported net income of $300,000 and $325,000 for 2017 and 2018, respectively. Pepe uses the equity method to account for its investment in Devin. Assuming there are no excess amortizations or other intra-entity transactions, compute the net income attributable to the noncontrolling interest of Devin for 2018. Multiple Choice $126,800. $130,000. $122,000. $130,800. $129,600.
Answer #1

Loss on sale of equipment =45000-(120000-66000) =9000

Loss amortization per year =9000/9 =1000

Net income of 2018 =325000

Income after loss recognition =324000

Net income attributable to the noncontrolling interest=324000*(1-60%) =129600